OpenStudy (anonymous):

(LOOK AT RESPONSE FOR SERIES) What is the conmmon difference of the series below?

OpenStudy (anonymous):

\[\sum_{n=2}^{6}2n+3\]

OpenStudy (anonymous):

(2*3+3)-(2*2+3)= (6+3)-(4+3)=9-7 =2

OpenStudy (anonymous):

2 is the common difference

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

my pleasure:)